### v1

# Extensions of the Busemann-Petty Problem for Arbitrary Measures

Abstract.
The classical Busemann–Petty problem asks whether smaller central hyperplane sections of origin-symmetric convex bodies necessarily imply smaller total volume. Zvavitch studied this question for arbitrary measures with continuous even densities, providing sufficient conditions for affirmative cases in terms of the distributional behavior of the ratio between the densities involved. We refine this result by extending it to a broader class of distributions and allowing a distinct pair of densities for each body—one for hyperplane sections and another for the full volume. We also present some examples illustrating cases not covered by previous results.
Key words and phrases:
Busemann–Petty problem, convex bodies, intersection bodies, positive-definite distributions
2020 Mathematics Subject Classification:
52A20, 52A21, 42B10, 44A12
1. Introduction
Understanding how lower-dimensional information determines global geometric quantities is a central theme in convex and geometric analysis. A prototypical example of this phenomenon is provided by the Busemann–Petty problem in $\mathbb{R}^{n}$, which compares (n1)(n-1)-dimensional central hyperplane sections of two origin-symmetric convex bodies and asks whether smaller sections necessarily imply smaller $n$-dimensional volume.
We recall that a convex body in $\mathbb{R}^{n}$ is a convex, compact set having the origin as an interior point. More generally, a star body is a subset $K\subseteq\mathbb{R}^{n}$ of the form
$K=\{x\in\mathbb{R}^{n}:x=0\text{ or }|x|\leq\rho(x/|x|)\},$
where $\rho:S^{n-1}\to(0,\infty)$ is continuous. The function $\rho=\rho_{K}$ is called the radial function of $K$. A body $K$ is said to be origin-symmetric (or simply symmetric) if K=KK=-K. Every convex body is, in particular, a star body. We may define the corresponding (quasi-)norm $\|x\|_{K}=\rho_{K}(x)^{-1}$, which is a norm when $K$ is a symmetric convex body.
The question posed by Busemann and Petty is as follows: Let $K,L\subseteq\mathbb{R}^{n}$ be symmetric convex bodies such that
$\left|K\cap H\right|_{n-1}\leq\left|L\cap H\right|_{n-1}$
for every hyperplane $H$ through the origin. Does it follow that $\left|K\right|_{n}\leq\left|L\right|_{n}$?
This interesting problem triggered a series of deep developments connecting convex geometry and Fourier analysis. The answer was shown to be affirmative for dimensions $n\leq 4$ and negative for $n\geq 5$, after major contributions culminating in the analytic characterization of intersection bodies due to Koldobsky (see, e.g., [BusemannPetty1956, Gardner1994, Koldobsky98b, Koldobsky98]).
A general version of the Busemann-Petty problem replaces the volumes of the sections and the bodies, by arbitrary measures with continuous densities. Consider a continuous even real function $f$ on $\mathbb{R}^{n}$. We define a measure $\mu$ whose density is $f$. For every hyperplane $H$,
(1) $\mu(K\cap H)=\int_{K\cap H}f(x)dx,$
is well-defined, where the integration is with respect to the (n1)(n-1) dimensional Lebesgue measure in $H$.
The Busemann-Petty problem for arbitrary measures was addressed by Zvavitch in [zvavitch2005busemann], where two different measures were considered, one for measuring the bodies, and a separate one for the sections. We state here the theorem by Zvavitch (see Section 2 below for the definition of positive-definiteness).
Theorem 1.
Let $f_{n-1}$ and $f_{n}$ be even continuous non-negative functions defined in $\mathbb{R}^{n}$ such that
(2) $\frac{f_{n-1}(tx)}{tf_{n}(tx)}$
is a non-increasing function of $t$ for any fixed $x\in S^{n-1}$. Suppose that a symmetric star body $K$ in $\mathbb{R}^{n}$ has the property that
(3) $\|x\|_{K}^{-1}\frac{f_{n}\left(\frac{x}{\|x\|_{K}}\right)}{f_{n-1}\left(\frac{x}{\|x\|_{K}}\right)}$
is a positive-definite distribution on $\mathbb{R}^{n}$.
Denote by $\mu_{n-1}$ and $\mu_{n}$ the measures associated with $f_{n-1}$ and $f_{n}$. Then for any symmetric star body $L$ in $\mathbb{R}^{n}$ satisfying
(4) $\mu_{n-1}(K\cap H)\leq\mu_{n-1}(L\cap H),$
for every hyperplane $H$ through the origin, we have
(5) $\mu_{n}(K)\leq\mu_{n}(L).$
This paper extends Zvavitch’s approach in two directions. First, we observe that a very mild modification in the argument allows the monotonicity assumption of the function in (2) to be reversed (so that a function previously required to be non-increasing may instead be taken non-decreasing), and, more generally, to be decomposed as a sum of a radially non-decreasing part and a radially non-increasing part. Second, through a modification of hypothesis (3), we permit different measurement schemes for each body: each may be measured with its own pair of densities (one for (n1)(n-1)-dimensional sections and one for $n$-dimensional volume). More precisely, $K$ is measured with the densities $f_{n-1}$ and $f_{n}$, while $L$ is measured with $g_{n-1}$ and $g_{n}$. It is worth pointing out that, in this direction, an isomorphic version of the Busemann–Petty problem with one measure assigned to each body was considered in [koldobsky2019measure, Theorem 1], where a comparison between measures was established in terms of the outer volume ratio distance to the class of generalized intersection bodies.
The two mentioned extensions lead to the following refinement of Theorem 1, obtained by modifying Zvavitch’s original argument.
Theorem 2.
Let $K,L\subseteq\mathbb{R}^{n}$ be origin-symmetric star bodies, and let $\mu_{n},\mu_{n-1},\nu_{n},\nu_{n-1}$ be four measures with non-negative even densities $f_{n},f_{n-1},g_{n},g_{n-1}$, where $f_{n-1},g_{n-1}$ are continuous. Define
(6) $h(x)=\frac{1}{|x|}\begin{cases}\frac{f_{n-1}(x)}{f_{n}(x)}&\text{if }x\in K\setminus L\\ \frac{g_{n-1}(x)}{g_{n}(x)}&\text{if }x\in L\setminus K\end{cases},$
and assume that there are non-negative functions $a,b:K\Delta L\to\mathbb{R}$ such that
(7) $h(x)=a(x)+b(x),$
where $a$ is radially non-decreasing and $b$ is radially non-increasing.
Assume that the function $F:\mathbb{R}^{n}\setminus\{0\}\to\mathbb{R}$ defined by
$F(x)=|x|^{-1}(a(x/\|x\|_{K})+b(x/\|x\|_{L}))^{-1}$
represents a positive-definite distribution.
If for every hyperplane $H$ passing through the origin,
(8) $\nu_{n-1}((L\setminus K)\cap H)\leq\mu_{n-1}((K\setminus L)\cap H),$
then
(9) $\nu_{n}(L\setminus K)\leq\mu_{n}(K\setminus L).$
Some important remarks are in order. Notice that Theorem 2 recovers Zvavitch’s result, Theorem 1. Indeed, when $\mu_{n-1}=\nu_{n-1}$ and $\mu_{n}=\nu_{n}$, condition (8) reduces to (4), since the contributions of the measures in $K\cap L$ cancel on both sides. With this formulation, inequality (9) makes sense even if the measures are not locally integrable at the origin. Moreover, although Zvavitch’s original statement required continuity of the density $f_{n}$, a careful inspection of the proof shows that this assumption may be omitted. Similarly, in our theorem, we do not require continuity of the densities $f_{n}$ and $g_{n}$ to measure the total mass of the bodies.
Finally we remark that, although the function in (2) does not correspond to Zvavitch’s original formulation of the theorem, where its reciprocal appears and a non-decreasing condition is required, we have chosen to present it this way to highlight the analogy with our generalization, in which this quotient can, in a certain sense, be expressed as a non-decreasing part and a non-increasing part.
The paper is organized as follows: in Section 2 we give all the necessary definitions, including positive-definite distributions, in Section 3 we give some examples of applications, and finally provide a proof of Theorem 2 in Section 4.
2. Preliminaries
We follow [GelfandVilenkin, KoldobskyBook] for standard definitions concerning the Fourier transform of distributions. We denote by $\mathcal{S}$ the Schwartz space of complex-valued test functions on $\mathbb{R}^{n}$—that is, $C^{\infty}$ functions decreasing faster than any polynomial together with all their derivatives.
A distribution on $\mathbb{R}^{n}$ is an element $f\in\mathcal{S}^{\prime}$, the dual space of $\mathcal{S}$. We denote by $\langle f,\varphi\rangle$ the action of $f\in\mathcal{S}^{\prime}$ on $\varphi\in\mathcal{S}$. Any locally integrable function $f$ defines a distribution by
$\varphi\in\mathcal{S}\ \mapsto\ \int_{\mathbb{R}^{n}}f(x)\,\varphi(x)\,dx,$
which is identified with $f$ itself. By abuse of notation, the right-hand side is also denoted $\langle f,\varphi\rangle$.
The Fourier transform of $\varphi\in\mathcal{S}$ is defined as
$\hat{\varphi}(\xi)=\int_{\mathbb{R}^{n}}\varphi(x)\,e^{-i(x,\xi)}\,dx,$
where $(\cdot,\cdot)$ denotes the scalar inner product of $\mathbb{R}^{n}$. For any distribution $f\in\mathcal{S}^{\prime}$, its Fourier transform $\hat{f}$ is defined by
$\langle\hat{f},\varphi\rangle=\langle f,\hat{\varphi}\rangle$.
If $p\in(-n,0)$ and $f$ is a homogeneous function of degree $p$, then $f$ is locally integrable and thus defines a distribution on $\mathcal{S}$. A distribution $f$ is said to be even homogeneous of degree $p$ if
$\langle f,\varphi(x/\lambda)\rangle=|\lambda|^{n+p}\langle f,\varphi\rangle$
for every $\varphi\in\mathcal{S}$ and every $\lambda\in\mathbb{R}\setminus\{0\}$. If $f$ is even homogeneous of degree $p$ and smooth, then $\hat{f}$ is even homogeneous of degree np-n-p and smooth as well.
A distribution $f$ is called positive if $\langle f,\varphi\rangle\geq 0$ for every nonnegative $\varphi\in\mathcal{S}$. It is called positive-definite if $\hat{f}$ is positive. Moreover, any positive-definite distribution is the Fourier transform of a nonnegative tempered measure (see [GelfandVilenkin, KoldobskyBook]).
Recall that a distribution $f\in\mathcal{S}^{\prime}$ is tempered if it defines a continuous linear functional on $\mathcal{S}$, i.e., if $\langle f,\varphi\rangle$ is well-defined and continuous for all $\varphi\in\mathcal{S}$. Equivalently, a (Borel) measure $\mu$ on $\mathbb{R}^{n}$ is tempered if it defines a tempered distribution, or, equivalently, if it has at most polynomial growth:
$\int_{\mathbb{R}^{n}}(1+|x|)^{-N}\,d\mu(x)<\infty$
for some integer $N\geq 0$. All finite measures and, in particular, all measures with compact support are tempered.
The Fourier transform of an even homogeneous distribution of degree $p\in(-n,0)$ is again an even homogeneous distribution of degree np-n-p (see [KoldobskyBook]).
Let $f$ be a homogeneous function of degree $n+p\in(-n,0)$ which is also a positive-definite distribution. Then the Fourier transform $\hat{f}$ is a tempered measure that can be expressed in polar coordinates (see [KoldobskyBook, Lemma 2.24]). Hence there exists a nonnegative measure $\gamma_{0}$ on $S^{n-1}$ such that
$\langle\hat{f},\varphi\rangle=\int_{S^{n-1}}\!d\gamma_{0}(\theta)\int_{0}^{\infty}r^{p-1}\varphi(r\theta)\,dr,\qquad\forall\,\varphi\in\mathcal{S}.$
The following proposition, known as the spherical Parseval formula, corresponds to [KoldobskyBook, Corollary 3.23] with $k=1$.
Proposition 3.
Let $f,g$ be two even functions on $\mathbb{R}^{n}$, homogeneous of degrees 1-1 and 1n1-n, respectively. Suppose that $f$ represents a positive-definite distribution on $\mathbb{R}^{n}$, with $\gamma_{0}$ the corresponding measure on $S^{n-1}$, and that $g\in C^{0}(\mathbb{R}^{n}\setminus\{0\})$. Then
$\int_{S^{n-1}}\hat{g}(v)\,d\gamma_{0}(v)=(2\pi)^{n}\int_{S^{n-1}}g(v)\,f(v)\,dv.$
We define the extension operator $E_{p}:C(S^{n-1})\to C(\mathbb{R}^{n}\setminus\{0\})$, for $p\in(-n,0)$, by
$E_{p}g(rv)=r^{p}g(v),\qquad r>0,\ v\in S^{n-1}.$
We also write $E_{p}^{\wedge}f=(E_{p}f)^{\wedge}$.
With this notation, the Parseval formula on the sphere, for $f,g\in C^{\infty}(S^{n-1})$, reads
$\int_{S^{n-1}}E_{k}^{\wedge}g(v)\,f(v)\,dv=\int_{S^{n-1}}g(v)\,E_{k}^{\wedge}f(v)\,dv.$
The Radon transform of a continuous function $f:S^{n-1}\to\mathbb{R}$ is defined by
$\mathcal{R}f(v)=\int_{S^{n-1}\cap v^{\perp}}f(w)\,dw,$
where $v^{\perp}$ is the subspace orthogonal to $v$, and integration is with respect to the (n2)(n-2)-dimensional Lebesgue measure on the subsphere.
The relation between the Fourier and Radon transforms is given by (see [KoldobskyBook, Lemma 3.7])
(10) $\mathcal{R}f=\frac{1}{\pi}\,E_{1-n}^{\wedge}f,$
for any $f\in C(S^{n-1})$.
Definition 4.
A star body $K$ is said to be an intersection body of a star body $L$ if, for every $v\in S^{n-1}$,
$\rho_{K}(v)=\left|L\cap v^{\perp}\right|_{n-1}=\frac{1}{n-1}\mathcal{R}(\rho_{L}^{n-1})(v).$
We say that $K$ is an intersection body if it is the limit, in the Hausdorff metric, of intersection bodies of star bodies.
Thanks to (10), Koldobsky obtained the following analytic characterization of intersection bodies.
Theorem 5 ([Koldobsky98, Theorem 1]).
An origin-symmetric star body $K\subseteq\mathbb{R}^{n}$ is an intersection body if and only if the Fourier transform of $\rho_{K}=\|\cdot\|_{K}^{-1}$ is a positive-definite distribution.
By Theorem 5, the hypothesis (3) is equivalent to the statement that the star body $M$ defined by
$\rho_{M}(x)=\|x\|_{K}^{-1}\frac{f_{n}\!\left(\tfrac{x}{\|x\|_{K}}\right)}{f_{n-1}\!\left(\tfrac{x}{\|x\|_{K}}\right)},\qquad x\in\mathbb{R}^{n}\setminus\{0\},$
is an intersection body.
Before proving Theorem 2, we present some examples.
3. Examples
We now present some cases not covered by Theorem 1. In the next example, we have only two measures $f_{n}=g_{n}$ and $f_{n-1}=g_{n-1}$ (same situation as in Theorem 1), but the function in (2) is non-decreasing instead of non-increasing.
Example 6.
Let $K,L\subseteq\mathbb{R}^{n}$ be star bodies such that
$\rho_{K}(v)^{-1}=\int_{M\cap v^{\perp}}|x|^{2}dx$
for some star body $M$. If for every $v\in S^{n-1}$,
$\int_{L\cap v^{\perp}}|x|^{2}dx\leq\int_{K\cap v^{\perp}}|x|^{2}dx,$
then $|L|_{n}\leq|K|_{n}$.
Proof.
Let $f_{n-1}=g_{n-1}=|x|^{2}$ and $f_{n}=g_{n}=1$. As in Theorem 2, we denote the corresponding measures by $\mu_{n-1},\nu_{n-1}$, $\mu_{n},\nu_{n}$ and continue to use the notation $h(x)$ from (7). Then we can write
(11) $h(x)=|x|=a(x)+b(x),$
where $a(x)=|x|$ and $b(x)=0$. By polar coordinates, for every $v\in S^{n-1}$,
(12) $\rho_{K}(v)^{-1}$ $=\int_{M\cap v^{\perp}}|x|^{2}dx$
(13) $=\frac{1}{n+1}\mathcal{R}(\rho_{M}^{n+1})(v),$
so that the star body $\bar{K}$ defined by $\rho_{\bar{K}}(v)=\rho_{K}(v)^{-1}$ for every $v\in S^{n-1}$, is the intersection body of the star body whose radial function coincides with $\rho_{M}^{\frac{n+1}{n-1}}$ in the unit sphere. By Theorem 5 the function
$F(x)=|x|^{-1}(|x/\|x\|_{K}|)^{-1}=\rho_{\bar{K}}(x)$
is positive definite and we can apply Theorem 2 to conclude that $|L|_{n}\leq|K|_{n}$. ∎
We stress that our result applies to total measures that may fail to be locally integrable at the origin, a condition required in the original hypotheses of Theorem 1. The following example illustrates this situation with the densities $f_{n}$ and $g_{n}$ taken as $1/|x|^{n}$.
Example 7.
Let $K,L\subseteq\mathbb{R}^{n}$ be star bodies such that the radial body $M$ defined by $\rho_{M}(v)=\rho_{K}^{n-1}(v)$ for $v\in S^{n-1}$, is an intersection body. If for every $v\in S^{n-1}$
(14) $|L\cap v^{\perp}|_{n-1}\leq|K\cap v^{\perp}|_{n-1}$
then
(15) $\int_{L\setminus K}|x|^{-n}dx\leq\int_{K\setminus L}|x|^{-n}dx.$
Proof.
Let $f_{n-1}=g_{n-1}=1$ and $f_{n}=g_{n}=|x|^{n}$. Then
(16) $q(x)=\frac{f_{n-1}(x)}{|x|f_{n}(x)}=|x|^{n-1},$
which is radially non-decreasing. Since, by hypothesis, the map $v\mapsto\rho_{K}^{\,n-1}(v)$ is positive-definite, Theorem 2 implies the claimed conclusion. ∎
Strictly speaking, not requiring local integrability does not constitute a genuine improvement, since the same conclusion can essentially be recovered from Theorem 1 by suitably modifying the measures in a small neighborhood of the origin. Nevertheless, our formulation allows one to interpret the theorem directly in terms of the total measures of the respective differences $K\setminus L$ and $L\setminus K$, without any assumptions about the behavior at the origin.
Let us now show an example for which the function $F$ is neither non-decreasing nor non-increasing. The example is admittedly artificial, but shows how the limitations of Theorem 1 are only technical.
Take
(17) $f_{n-1}(x)=\varepsilon|x|^{3}+|x|,$ $f_{n}(x)=|x|$
(18) $g_{n-1}(x)=\varepsilon|x|^{5}+|x|^{3},$ $g_{n}(x)=|x|^{3},$
where $\varepsilon>0$ is small. then
$F(x)=\varepsilon|x|+\frac{1}{|x|},$
and we may take
$a(x)=\varepsilon|x|,\qquad b(x)=\frac{1}{|x|}.$
$F(x)=|x|^{-1}(\varepsilon|x/\|x\|_{K}|+|x/\|x\|_{L}|^{-1})^{-1}$
We obtain the following:
Example 8.
Assume $K,L$ are radial bodies such that the radial body $M$ defined by
$\rho_{M}(v)=\left(\varepsilon\rho_{K}(v)+\frac{1}{\rho_{L}(v)}\right)^{-1},\quad v\in S^{n-1},$
is an intersection body.
If for every $\xi\in S^{n-1}$
$\int_{(L\setminus K)\cap\xi^{\perp}}(\varepsilon|x|^{5}+|x|^{3})dx\leq\int_{(K\setminus L)\cap\xi^{\perp}}(\varepsilon|x|^{3}+|x|)dx,$
then
$\int_{L\setminus K}|x|^{3}\leq\int_{K\setminus L}|x|.$
If $K,L$ are $C^{2}$-smooth convex bodies and $L$ has strictly-positive Gauss curvature, then $M$ will be convex for small $\varepsilon$. The hypothesis on $M$ is thus satisfied automatically in dimension $n\leq 4$, by the fact that every convex body in dimension $n\leq 4$ is an intersection body.
4. Proof of Theorem 2
Before proceeding to the proof of our main theorem, we establish a simple but useful continuity result that will be needed in what follows.
Lemma 9.
Let $f,g:\mathbb{R}^{n}\to\mathbb{R}$ be two continuous functions, and $K,L$ two star bodies. Define $Q:S^{n-1}\to\mathbb{R}$ by
$Q(v)=\int_{\rho_{L}(v)}^{\rho_{K}(v)}q(rv)dr,$
where
(19) $q(x)=\begin{cases}f(x)&\text{if }x\in K\setminus L\\ g(x)&\text{if }x\in L\setminus K\\ 0&\text{otherwise.}\end{cases}$
Then $Q$ is a continuous function.
Proof.
Let $v_{k}\in S^{n-1}$ be a sequence converging to $v\in S^{n-1}$ as $k\to\infty$. We distinguish two separate cases, $\rho_{K}(v)=\rho_{L}(v)$ and $\rho_{K}(v)\neq\rho_{L}(v)$. In the first case we have $Q(v)=0$. Since $q$ is a bounded function in the compact set $K\cup L$,
$|Q(v_{k})|\leq C|\rho_{K}(v_{k})-\rho_{L}(v_{k})|\to C|\rho_{K}(v)-\rho_{L}(v)|=0$
as $k\to\infty$.
In the second case we may assume without loss of generality that $\rho_{L}(v)<\rho_{K}(v)$. By the continuity of $\rho_{L},\rho_{K}$ we have $\rho_{L}(v_{k})<\rho_{K}(v_{k})$ for $k$ sufficiently large. Moreover, for any $r\in(\rho_{L}(v),\rho_{K}(v))$ we have $\rho_{L}(v_{k})<r<\rho_{K}(v_{k})$ (for $k$ sufficiently large). For any such $r$ we have $q(rv_{k})=f(rv_{k})$ which converges pointwise to $f(rv)=q(rv)$ as $k\to\infty$, by the continuity of $f$. Then
$q(rv_{k})1_{(\rho_{L}(v_{k}),\rho_{K}(v_{k}))}(r)\to q(rv)1_{(\rho_{L}(v),\rho_{K}(v))}(r)$
pointwise in $(\rho_{L}(v),\rho_{K}(v))$. Now we may use dominated convergence to obtain
$Q(v_{k})=\int_{\rho_{L}(v_{k})}^{\rho_{K}(v_{k})}f(rv_{k})dr\to\int_{\rho_{L}(v)}^{\rho_{K}(v)}q(rv)dr=Q(v),$
and the lemma is proven. ∎
We now have all the ingredients needed to prove our main theorem.
Proof of Theorem 2.
Define for i=n1,ni=n-1,n,
$h_{i}(x)=\begin{cases}f_{i}(x)&x\in K\setminus L\\ g_{i}(x)&x\in L\setminus K\end{cases},$
then the hypothesis
$\nu_{n-1}((L\setminus K)\cap\xi^{\perp})\leq\mu_{n-1}((K\setminus L)\cap\xi^{\perp})$
for every $\xi\in S^{n-1}$ is equivalent to
$\int_{(K\setminus L)\cap\xi^{\perp}}h_{n-1}(x)dx-\int_{(L\setminus K)\cap\xi^{\perp}}h_{n-1}(x)dx\geq 0.$
Using polar coordinates and relation (10), we write this as
(20) $0$ $\leq\int_{S^{n-1}\cap\xi^{\perp}}\int_{\rho_{L}(v)}^{\rho_{K}(v)}r^{n-2}h_{n-1}(rv)drdv$
(21) $=\mathcal{R}_{v}\left[\int_{\rho_{L}(v)}^{\rho_{K}(v)}r^{n-2}h_{n-1}(rv)dr\right](\xi)$
(22) $=\frac{1}{\pi}(E_{1-n}^{\wedge})_{v}\left[\int_{\rho_{L}(v)}^{\rho_{K}(v)}r^{n-2}h_{n-1}(rv)dr\right](\xi).$
Here $\mathcal{R}_{v},(E_{1-n}^{\wedge})_{v}$ are the operators defined previously, with respect to the variable $v$.
Let $\gamma_{0}$ the tempered measure corresponding to the positive definite distribution $|x|^{-1}(a(x/\|x\|_{K})+b(x/\|x\|_{L}))^{-1}$. Integrating with respect to $\gamma_{0}$ we obtain
(23) $0\leq\int_{S^{n-1}}(E_{1-n}^{\wedge})_{v}\left[\int_{\rho_{L}(v)}^{\rho_{K}(v)}r^{n-2}h_{n-1}(rv)dr\right](\xi)d\gamma_{0}(\xi).$
By Lemma 9 applied to the functions $|x|^{n-2}f_{n-1}(x)$ and $|x|^{n-2}g_{n-1}(x)$, the mapping $\left[v\mapsto\int_{\rho_{L}(v)}^{\rho_{K}(v)}r^{n-2}h_{n-1}(rv)dr\right]$ is continuous in $S^{n-1}$. Thus, by Proposition 3 and equations (6) and (7), inequality (23) becomes
(24) $0\leq\int_{S^{n-1}}(a(v/\|v\|_{K})+b(v/\|v\|_{L}))^{-1}\int_{\rho_{L}(v)}^{\rho_{K}(v)}r^{n-2}h_{n-1}(rv)drdv\\ =\int_{S^{n-1}}(a(v/\|v\|_{K})+b(v/\|v\|_{L}))^{-1}\int_{\rho_{L}(v)}^{\rho_{K}(v)}r^{n-1}h_{n}(rv)(a(rv)+b(rv))drdv.$
Now, since $a(rv)$ is non-decreasing and $b(rv)$ is non-increasing with respect to $r$, we can bound
$\int_{\rho_{L}(v)}^{\rho_{K}(v)}r^{n-1}h_{n}(rv)(a(rv)+b(rv))dr\leq(a(\rho_{K}(v)v)+b(\rho_{L}(v)v))\int_{\rho_{L}(v)}^{\rho_{K}(v)}r^{n-1}h_{n}(rv)dr.$
Combining this with (24) we obtain
(25) $0$ $\leq\int_{S^{n-1}}\int_{\rho_{L}(v)}^{\rho_{K}(v)}r^{n-1}h_{n}(rv)drdv$
(26) $=\int_{K\Delta L}h_{n}(x)dx=\mu_{n}(K\setminus L)-\nu_{n}(L\setminus K),$
and the proof is complete. ∎

### v2

@@
 This interesting problem triggered a series of deep developments connecting convex geometry and Fourier analysis.
-The answer was shown to be affirmative for dimensions $n\leq 4$ and negative for $n\geq 5$, after major contributions culminating in the analytic characterization of intersection bodies due to Koldobsky (see, e.g., [BusemannPetty1956, Gardner1994, Koldobsky98b, Koldobsky98]).
+The answer was shown to be affirmative for dimensions $n\leq 4$ and negative for $n\geq 5$, after major contributions culminating in the analytic characterization of intersection bodies.
+See [1, 2, 4, 5, 6] for the solution of the Busemann-Petty problem, and its history.
 A general version of the Busemann-Petty problem replaces the volumes of the sections and the bodies, by arbitrary measures with continuous densities.
@@
 is well-defined, where the integration is with respect to the (n1)(n-1) dimensional Lebesgue measure in $H$.
-The Busemann-Petty problem for arbitrary measures was addressed by Zvavitch in [zvavitch2005busemann], where two different measures were considered, one for measuring the bodies, and a separate one for the sections.
+The Busemann-Petty problem for arbitrary measures was addressed by Zvavitch in [8], where two different measures were considered, one for measuring the bodies, and a separate one for the sections.
 We state here the theorem by Zvavitch (see Section 2 below for the definition of positive-definiteness).
@@
 More precisely, $K$ is measured with the densities $f_{n-1}$ and $f_{n}$, while $L$ is measured with $g_{n-1}$ and $g_{n}$.
-It is worth pointing out that, in this direction, an isomorphic version of the Busemann–Petty problem with one measure assigned to each body was considered in [koldobsky2019measure, Theorem 1], where a comparison between measures was established in terms of the outer volume ratio distance to the class of generalized intersection bodies.
+It is worth pointing out that, in this direction, an isomorphic version of the Busemann–Petty problem with one measure assigned to each body was considered in [7, Theorem 1], where a comparison between measures was established in terms of the outer volume ratio distance to the class of generalized intersection bodies.
 The two mentioned extensions lead to the following refinement of Theorem 1, obtained by modifying Zvavitch’s original argument.
@@
 2. Preliminaries
-We follow [GelfandVilenkin, KoldobskyBook] for standard definitions concerning the Fourier transform of distributions.
+We follow [3, 6] for standard definitions concerning the Fourier transform of distributions.
 We denote by $\mathcal{S}$ the Schwartz space of complex-valued test functions on $\mathbb{R}^{n}$—that is, $C^{\infty}$ functions decreasing faster than any polynomial together with all their derivatives.
@@
 It is called positive-definite if $\hat{f}$ is positive.
-Moreover, any positive-definite distribution is the Fourier transform of a nonnegative tempered measure (see [GelfandVilenkin, KoldobskyBook]).
+Moreover, any positive-definite distribution is the Fourier transform of a nonnegative tempered measure (see [3, 6]).
 Recall that a distribution $f\in\mathcal{S}^{\prime}$ is tempered if it defines a continuous linear functional on $\mathcal{S}$, i.e., if $\langle f,\varphi\rangle$ is well-defined and continuous for all $\varphi\in\mathcal{S}$.
@@
 for some integer $N\geq 0$. All finite measures and, in particular, all measures with compact support are tempered.
-The Fourier transform of an even homogeneous distribution of degree $p\in(-n,0)$ is again an even homogeneous distribution of degree np-n-p (see [KoldobskyBook]).
+The Fourier transform of an even homogeneous distribution of degree $p\in(-n,0)$ is again an even homogeneous distribution of degree np-n-p (see [6]).
 Let $f$ be a homogeneous function of degree $n+p\in(-n,0)$ which is also a positive-definite distribution.
-Then the Fourier transform $\hat{f}$ is a tempered measure that can be expressed in polar coordinates (see [KoldobskyBook, Lemma 2.24]).
+Then the Fourier transform $\hat{f}$ is a tempered measure that can be expressed in polar coordinates (see [6, Lemma 2.24]).
 Hence there exists a nonnegative measure $\gamma_{0}$ on $S^{n-1}$ such that
 $\langle\hat{f},\varphi\rangle=\int_{S^{n-1}}\!d\gamma_{0}(\theta)\int_{0}^{\infty}r^{p-1}\varphi(r\theta)\,dr,\qquad\forall\,\varphi\in\mathcal{S}.$
-The following proposition, known as the spherical Parseval formula, corresponds to [KoldobskyBook, Corollary 3.23] with $k=1$.
+The following proposition, known as the spherical Parseval formula, corresponds to [6, Corollary 3.23] with $k=1$.
 Proposition 3.
@@
 where $v^{\perp}$ is the subspace orthogonal to $v$, and integration is with respect to the (n2)(n-2)-dimensional Lebesgue measure on the subsphere.
-The relation between the Fourier and Radon transforms is given by (see [KoldobskyBook, Lemma 3.7])
+The relation between the Fourier and Radon transforms is given by (see [6, Lemma 3.7])
 (10) $\mathcal{R}f=\frac{1}{\pi}\,E_{1-n}^{\wedge}f,$
@@
 Thanks to (10), Koldobsky obtained the following analytic characterization of intersection bodies.
-Theorem 5 ([Koldobsky98, Theorem 1]).
+Theorem 5 ([5, Theorem 1]).
 An origin-symmetric star body $K\subseteq\mathbb{R}^{n}$ is an intersection body if and only if the Fourier transform of $\rho_{K}=\|\cdot\|_{K}^{-1}$ is a positive-definite distribution.